A sulfuric acid solution containing 571.8 gg of H2SO4H2SO4 per liter of aqueous solution has a density of 1.329 g/cm3g/cm3.

Part A

Calculate the mass percentage of H2SO4H2SO4 in this solution.

Express your answer using four significant figures.

Part B

Calculate the mole fraction of H2SO4H2SO4 in this solution.

Express your answer using four significant figures.

Part C

Calculate the molality of H2SO4H2SO4 in this solution.

Express your answer using four significant figures.

Part D

Calculate the molarity of H2SO4H2SO4 in this solution.

Express your answer using four significant figures.
Answer #1

density is 1.329 g/ cm3 I.e 1329 grams in 1 liter or 1000mL and the mass of the H2SO4 is 571.8 g

mass percent =  mass of the H2SO4 / total mass * 100

mass percent = 571.8 / 1329 * 100 => 43.0248 % => 43.02%

moles = mass / molar mass

moles = 571.8 / 98 => 5.8347 moles

no of moles of H20 => 1329 - 571.8 => 757.2 grams

moles => 757.2 g / 18 g/mol => 42.066 moles

mole fraction = 5.8347 / (5.8347 + 42.066) => 0.1218  

molality => 5.8347 mol / 0.7572 kg => 7.706 m

molarity => 5.8347 mol / 1 liter => 5.835 M